OpenStudy (anonymous):

\[ \lim_{n=1, 2, \1dots} a_n, \qquad \max_{x

OpenStudy (anonymous):

\[ \lim_{n=1, 2, \dots} a_n, \qquad \max_{x<X} x \]

OpenStudy (anonymous):

\[ \frac{n(n+1)}{2}, \quad \frac{\frac{\sqrt{x}+1}{2}-x}{y^2} \]

OpenStudy (anonymous):

\sqrt[64]{x} = \sqrt{\sqrt{\sqrt{\sqrt{\sqrt{\sqrt{x}}}}}}

OpenStudy (unklerhaukus):

\[\sqrt[64]{x} = \sqrt{\sqrt{\sqrt{\sqrt{\sqrt{\sqrt{x}}}}}}\] ... radical!

OpenStudy (anonymous):

Haha, I didn't know properly the radical code. so what was the problem? @UnkleRhaukus

OpenStudy (unklerhaukus):

you got the code right, just missing the delimiters\[\[\]\[\]\]